# Robust Hyperbolic Learning with Curvature-Aware Optimization

### v3

@@ 3.3 Riemannian AdamW Optimizer
The AdamW optimizer was first introduced by Loshchilov & Hutter (2019) and relies on an improved application of the L2-regularization factor in the base Adam optimizer. L2-regularization works on the principle that networks with smaller weights tend to have better generalization performance than equal networks with higher weight values. Adam applied the L2-regularization during the gradient update step by incorporating it in the loss. However, Loshchilov & Hutter (2019) argue that this is inconsistent since the regularization effect is reduced by the magnitude of the gradient norms. Instead, they apply the weight decay directly during the parameter update step. AdamW is then shown to generalize better and lead to better convergence and has become a popular choice for many vision tasks.
Given the above, we believe AdamW is significant for hyperbolic learning. Hyperbolic spaces, with their higher representational capacity, are prone to overfitting, especially under data scarcity during training (Gao et al., 2022). Thus, AdamW’s enhanced L2-regularization could improve hyperbolic models.
@@
In the following, we derive AdamW for the Lorentz manifold and suggest its extension to the Poincaré ball. The key difference between AdamW and Riemannian Adam lies in direct weight regularization, which is challenging in Lorentz space due to the lack of an intuitive subtraction operation. To address this, we re-frame parameter regularization in AdamW as a weighted centroid with the origin $\overline{\bm{0}}$
$\bm{\theta_{t-1}}-\gamma\lambda\bm{\theta_{t-1}}=(1-\gamma\lambda)\bm{\theta_{t-1}}+\gamma\lambda\overline{\bm{0}}$
where $\gamma$ is the learning rate and $\lambda$ is the weight decay value. The centroid operation is well defined in the Lorentz space which now allows for a direct translation of AdamW. The regularization schema becomes:
$\bm{\theta_{t}}=\begin{dcases}\mu_{\mathbb{L}}([\bm{\theta_{t-1}},\overline{\bm{0}}],\bm{\nu}=[1-\gamma\lambda,\gamma\lambda])&\quad\textrm{if }\bm{\theta}\in\mathbb{L}\\ \bm{\theta_{t-1}}-\gamma\bm{\theta_{t-1}}\lambda&\quad\textrm{otherwise}\end{dcases}$
where $\mu_{\mathbb{L}}$ is the Lorentz centroid, and $\bm{\nu}$ are the weights. By removing the later gradient decay and introducing this operation, we adapt AdamW for use in the Lorentz space.
Algorithm 1 Curvature Learning Aware Optimization
  Given parameters $\bm{\theta}=[\bm{\theta}_{euclid},\bm{\theta}_{\mathbb{L}},\bm{\theta}_{K}]$ and gradients denoted $\mathcal{G}$
  function Optimizer Step
     for $\bm{p}$ in $\bm{\theta}_{\mathbb{L}}$ do
        Vanilla Riemannian Optimizer Step
     end for
     for $\bm{p}$ in $\bm{\theta}_{euclid}\cup\bm{\theta}_{K}$ do
        Vanilla Euclidean Optimizer Step
     end for
     MoveParameters()
  end function
  function Move Parameters
     for $\bm{p}$ in $\bm{\theta}_{\mathbb{L}}$ do
        $\bm{\mathcal{G}}_{temp}=\mathrm{PT}^{K_{t-1}}_{{\bm{p}}\rightarrow{\bm{\overline{\bm{0}}}_{t-1}}}\left({\bm{\mathcal{G}}}\right)$
        $\bm{z}=\log_{\bm{\overline{\bm{0}}}_{t-1}}^{K_{t-1}}(\bm{p})$
        $\bm{p}=\exp_{\overline{\bm{0}}_{t}}^{K_{t}}(\bm{z})$
        $\bm{\mathcal{G}}=\mathrm{PT}^{K_{t}}_{{\bm{\overline{\bm{0}}}_{t}}\rightarrow{\bm{p}}}\left({\bm{\mathcal{G}}_{temp}}\right)$
     end for
  end function
@@ 3.4 Maximum Distance Rescaling
$D^{K}_{\overline{\bm{0}}_{max}}=arccosh(\frac{x_{t_{max}}}{\sqrt{K}})\cdot\sqrt{K}$ (2)
Under Float32 precision, and to account for values of $K<1$ we use a limit value of $x_{t_{max}}=2\cdot 10^{3}$. When projected onto the tangent space of the origin, and with $K=1$, this translates to $||\log^{1}_{\overline{\bm{0}}}(\bm{x})||=D^{1}_{\overline{\bm{0}}_{max}}=9.1$. This value changes considerably as the value of $K$ changes. Vectors outside this radius lead to instability and performance degradation due to inaccurate approximation. This problem is only exacerbated as the dimensionality of the hyperbolic vector increases. Higher dimensional vectors tend to have larger norms which limits hyperbolic models’ abilities to scale up.
To constrain hyperbolic vectors within a specified maximum distance, either a normalization function or a parameter clipping method is required. Parameter clipping can be challenging to train, as it may lead to information loss and introduce non-smooth gradients. On the other hand, common normalization functions like tanh and the sigmoid function tend to saturate quickly, limiting their effectiveness as seen in the sigmoid implementation by Chen et al. (2022).
@@
To address these issues, we introduce a modified scaling function, designed to provide finer control over both the maximum values and the slope of the curve. A visualization of this function is provided in Figure 2, and the formulation is presented below:
$\bm{y}_{rescaled}=\frac{\bm{y}}{\lVert\bm{y}\rVert}\cdot m\cdot tanh(\lVert\bm{y}\rVert\cdot\frac{atanh(0.99)}{s\cdot m})$ (3)
where $y\in\mathbb{R}^{d}$, $m$ is our desired maximum value, and $s$ controls the slope of the curve. We now have a maximum distance value to adhere to and a flexible distance normalizing function. To apply this to the hyperbolic vectors, we suggest performing the scaling on the tangent plane of the origin. However, this is an expensive operation to perform repeatedly, as such we derive in Appendix C the equivalent factorized form for the scaling of the space values:
where $D(\bm{x},\overline{\bm{0}})_{rescaled}^{k}$ is the scaled distances by plugging in $D(\bm{x},\overline{\bm{0}})$ and $D^{K}_{\overline{\bm{0}}_{max}}$ in Equation 3. The time component of the vector is recalculated based on the norm of its spatial components, providing a complete scaling operation for the Lorentz space. We apply this tanh scaling during transitions between manifolds, including from Euclidean to Lorentz spaces and between Lorentz spaces of different curvatures. It is also used after Lorentz Boosts and direct Lorentz concatenations (Qu & Zou, 2022), as well as following variance-based rescaling in the batch normalization layer, where adjustments can push points outside the valid radius, often resulting in invalid values.
4 Domain Adaption
In order to empirically prove the effectiveness of our proposed solutions, we apply them to metric learning, EEG classification, image classification, and image generation tasks. In the following, we tackle each of these tasks by offering a brief overview of the problem setting with an existing reference model, our hyperbolic adaptation for this problem, and the goal from this experimental setup.
@@ 4.1 Hierarchical Metric Learning Problem
In their paper Kim et al. (2023) take on the problem of hierarchical clustering using an unsupervised hyperbolic loss regularizer they name HIER. This method relies on the use of hierarchical proxies as learnable ancestors of the embedded data points in hyperbolic space. Given a triplet of points ${xi,xj,xk}$ where $x_{i}$ and $x_{j}$ are determined to be related by a reciprocal nearest neighbor measure, and $x_{k}$ is an unrelated point, the HIER loss regularizer is then calculated as
where $D_{B}$ denotes the hyperbolic distance on the Poincaré ball, and $\rho_{ij}$ is the most likely least common ancestor of points $x_{i}$ and $x_{j}$. This encourages a smaller hyperbolic distance between $x_{i}$, $x_{j}$, and $\rho_{ij}$, and a larger distance with $\rho_{ijk}$. The opposite signal is then applied in the case of $x_{k}$, the irrelevant data point. Kim et al. (2023) show substantial performance uplifts for the HIER loss when applied to a variety of network architectures.
In the following experiment, we extend HIER to the Lorentz model (LHIER) and compare against the results provided by Kim et al. (2023).
@@
To adapt the HIER model to the hyperboloid, we first replace the Euclidean layer norm and linear layer with a Lorentzian norm and linear layer (Bdeir et al., 2024). We then modify the HIER loss by setting the hierarchical proxies as learnable hyperbolic parameters and replace the Poincaré distance with the Lorentzain distance. This allows us to optimize the hyperbolic parameters directly on the manifold using our new optimization schema. We also apply distance rescaling before layer norm and after the linear layer. We refer to this setup as LHIER.
@@
Kim et al. (2023) employ the Euclidean AdamW optimizer with Euclidean parameterizations of hyperbolic proxies. Their experiments use FP16 precision. As the setting is already hyperbolic, significant gains from transitioning to Lorentz space are unlikely. We use this setup to evaluate our components’ ability to learn curvature in low-precision, unstable environments. Any improvements likely stem from better curvature adaptation to the data and task.
Experimental Setup
We follow the experimental setup in Kim et al. (2023) and rely on four main datasets: CUB-200-2011 (CUB)(Welinder et al., 2010), Cars-196 (Cars)(Krause et al., 2013), Stanford Online Product (SOP)(Song et al., 2016), and In-shop Clothes Retrieval (InShop)(Liu et al., 2016). Performance is measured using Recall@k, the fraction of queries with at least one relevant sample in their k-nearest neighbors. All model backbones are pre-trained on ImageNet for fair comparisons.
@@ Results
Table 1: Performance of metric learning methods on the four datasets as provided by (Kim et al., 2023). All architecture backbones are pretrained and tested with the new LHIER loss. Superscripts denote their embedding dimensions and $\dagger$ indicates models using larger input images. As in (Kim et al., 2023), network architectures are abbreviated as, R–ResNet50 (Wang et al., 2017), B–Inception with BatchNorm (Ioffe & Szegedy, 2015), De–DeiT (Touvron et al., 2021), DN–DINO (Caron et al., 2021) and V–ViT (Dosovitskiy et al., 2020)
@@
NSoftmax R128 56.5 69.6 81.6 88.7 75.2 88.7
MIC R128 66.1 76.8 82.6 89.1 77.2 89.4
XBM R128 - - - - 80.6 91.6
XBM B512 65.8 75.9 82.0 88.7 79.5 90.8
HTL B512 57.1 68.8 81.4 88.0 74.8 88.3
MS B512 65.7 77.0 84.1 90.4 78.2 90.5
SoftTriple B512 65.4 76.4 84.5 90.7 78.6 86.6
PA B512 68.4 79.2 86.1 91.7 79.1 90.8
NSoftmax R512 61.3 73.9 84.2 90.4 78.2 90.6
ProxyNCA++ R512 69.0 79.8 86.5 92.5 80.7 92.0
Hyp R512 65.5 76.2 81.9 88.8 79.9 91.5
HIER R512 70.1 79.4 88.2 93.0 80.2 91.5
@@
Increase % 4.71 3.78 2.04 1.08 1.49 1.20
As shown in Table 1, LHIER learns curvature without stability issues. Our approach improves model performance, with recall@1 gains ranging from 1.49% to 4.71%. Additional ablation experiments with different optimizers and configurations are detailed in Table 7.
@@ 4.2 EEG Classification Problem
Given EEG recordings labeled into distinct categories, the goal is to classify new recordings based on signal patterns. Each recording is a time series $X\in\mathbb{R}^{C\times T}$, where $C$ is the number of channels and $T$ is the time length. Each label is $y\in\{1,\dots,K\}$. Given $N$ labeled recordings $(X_{1},y_{1}),\dots,(X_{N},y_{N})$ from an unknown distribution $p$, the task is to train a model $\hat{y}$ that maps EEG signals $X$ to their correct class.
@@
The reference model MAtt (Pan et al., 2022) has already taken steps towards improving data preparation. In the feature extraction phase, MAtt uses 2 convolutions along the time and sensor dimensions respectively. The embeddings are transformed into Symmetric Positive Definite (SPD) data by mapping the feature embeddings from Euclidean space onto the SPD manifold. The resulting sequence is then fed into a Manifold Attention Model and the outputs are projected back to the Euclidean space and classified using a fully connected layer.
Moving to Lorentz Space
To adapt Matt to a Lorentz space setting, we project the embeddings from the SPD manifold to the Lorentz manifold after the input slicing and sample covariance matrix projection. We then pass them through an adapted Lorentz Attention layer that replaces the SPD layers, distance function, and centroid with the Lorentzian counterparts. Finally, we use the Lorentz MLR (Bdeir et al., 2024) as a final classification head. We refer to this enhanced model as HyperMAtt.
@@
While this process already represents EEG sequences in Riemannian space, we enhance it further with our Riemannian AdamW Optimizer. Hyperbolic spaces are prone to overfitting, especially in EEG classification, where models are trained per subject with limited sessions, leading to data scarcity. The proposed optimizer improves regularization, ensuring smoother and more balanced convergence. We avoid curvature learning, as it could worsen overfitting and reduce performance, but retain distance rescaling functions.

### v4

@@ 3.3 Riemannian AdamW Optimizer
The AdamW optimizer was first introduced by Loshchilov and Hutter [26] and relies on an improved application of the L2-regularization factor in the base Adam optimizer. L2-regularization works on the principle that networks with smaller weights tend to have better generalization performance than equal networks with higher weight values. Adam applied the L2-regularization during the gradient update step by incorporating it in the loss. However, Loshchilov and Hutter [26] argue that this is inconsistent since the regularization effect is reduced by the magnitude of the gradient norms. Instead, they apply the weight decay directly during the parameter update step. AdamW is then shown to generalize better and lead to better convergence and has become a popular choice for many vision tasks.
Given the above, we believe AdamW is significant for hyperbolic learning. Hyperbolic spaces, with their higher representational capacity, are prone to overfitting, especially under data scarcity during training [12]. Thus, AdamW’s enhanced L2-regularization could improve hyperbolic models.
@@
In the following, we derive AdamW for the Lorentz manifold and suggest its extension to the Poincaré ball. The key difference between AdamW and Riemannian Adam lies in direct weight regularization, which is challenging in Lorentz space due to the lack of an intuitive subtraction operation. One option would be following the exponential mapping and retraction typically used for the optimization step. However we propose a simpler operation that reduces the need for expensive and less stable parallel transport operations. Specifically, we re-frame parameter regularization in AdamW as a weighted centroid with the origin
$\bm{\theta_{t-1}}-\gamma\lambda\bm{\theta_{t-1}}=(1-\gamma\lambda)\bm{\theta_{t-1}}+\gamma\lambda\bm{O}$
where $\gamma$ is the learning rate and $\lambda$ is the weight decay value. We can now directly translate this for hyperbolic parameters $\bm{\theta}_{\mathbb{L}}$ as $\mu^{\bm{\nu}}_{\mathbb{L}}([\bm{\theta_{t-1}},\overline{\bm{0}}])$ where $\mu^{\bm{\nu}}_{\mathbb{L}}$ is the weighted Lorentz centroid defined in Law et al. [22] and described in Appendix A. We define the centroid weights as $\bm{\nu}=[1-\gamma\lambda,\gamma\lambda]$. By removing the later gradient decay and introducing this operation as seen in Algorithm 2, we adapt AdamW for use in the Lorentz space.
Algorithm 1 Tangent Based Manifold Mapping
1:Given:
2:Hyperbolic parameters $\bm{\theta}^{K}_{\mathbb{L}}$ on the $K$-curve manifold, Parameter gradients $\mathcal{G}\in\mathcal{T}_{\bm{\theta}^{K}_{\mathbb{L}}}$
3:function Map Parameters
4:  for each $\bm{p}\in\bm{\theta}_{\mathbb{L}}$ do
5:   $\mathcal{G}_{\text{temp}}\leftarrow\mathcal{T}_{\bm{p}\rightarrow\overline{\bm{0}}_{t-1}}(\mathcal{G})$ $\triangleright$ Parallel transport gradient to previous origin
6:   $\bm{z}\leftarrow\log^{K_{t-1}}_{\overline{\bm{0}}_{t-1}}(\bm{p})$ $\triangleright$ Project parameter onto tangent space
7:   $\bm{p}\leftarrow\exp^{K_{t}}_{\overline{\bm{0}}_{t}}(\bm{z})$ $\triangleright$ Project back onto updated manifold
8:   $\mathcal{G}\leftarrow\mathcal{T}_{\overline{\bm{0}}_{t}\rightarrow\bm{p}}(\mathcal{G}_{\text{temp}})$ $\triangleright$ Transport gradient back for next update
9:  end for
10:end function
Algorithm 2 Riemannian Adam (RAdam) and Riemannian AdamW (RAdamW)
1:Manifold $\mathcal{M}$, initial parameters $\bm{\theta}_{\mathbb{L}}\in\mathcal{M}$
2:Learning rate $\alpha>0$, weight decay $\lambda\geq 0$, exponential decay rates $\beta_{1},\beta_{2}\in[0,1)$
[…]
4:Small constant $\epsilon>0$, max iterations $T$
5:Optimized parameters $p_{t}\in\bm{\theta}_{\mathbb{L}}$
6:Initialize moment vectors $m_{0}\leftarrow 0\in T_{p_{0}}\mathcal{M}$, $v_{0}\leftarrow 0\in T_{p_{0}}\mathcal{M}$
7:Initialize timestep $t\leftarrow 0$
8:for $t=1$ to $T$ do
9:  $g_{t}\leftarrow\mathrm{grad}~f(p_{t-1})$ $\lambda\cdot p_{t-1}$
10:  $\bm{\nu}\leftarrow[1-\gamma\lambda,\gamma\lambda]$
11:  $p_{t-1}\leftarrow\mu^{\bm{\nu}}_{\mathbb{L}}([\bm{p_{t-1}},\overline{\bm{0}}])$
12:  $g_{t}\leftarrow\text{egrad2rgrad}_{p_{t-1}}(g_{t})$ $\triangleright$ See Appendix A
13:  $m_{t}\leftarrow\beta_{1}\cdot m_{t-1}+(1-\beta_{1})\cdot g_{t}$
14:  $v_{t}\leftarrow\beta_{2}\cdot v_{t-1}+(1-\beta_{2})\cdot g_{t}\odot g_{t}$
15:  $\hat{m}_{t}\leftarrow m_{t}/(1-\beta_{1}^{t})$
16:  $\hat{v}_{t}\leftarrow v_{t}/(1-\beta_{2}^{t})$
17:  $\eta_{t}\leftarrow\alpha\cdot\frac{\hat{m}_{t}}{\sqrt{\hat{v}_{t}}+\epsilon}$
18:  $p_{t}\leftarrow\text{Retract}_{p_{t-1}}(-\eta_{t})$ $\triangleright$ See Appendix A
19:  $m_{t}\leftarrow\mathcal{T}_{p_{t-1}\to p_{t}}(m_{t})$
20:end forreturn $p_{t}$
@@ 3.4 Maximum Distance Rescaling
$D^{K}_{\overline{\bm{0}}_{\text{max}}}=\text{acosh}\left(\frac{x_{t_{max}}}{\sqrt{K}}\right)\cdot\sqrt{K}$ (1)
Under Float32 precision, and to account for values of $K<1$ we use a limit value of $x_{t_{max}}=2\cdot 10^{3}$. When projected onto the tangent space of the origin, this translates to $\text{log}_{\overline{\bm{0}}}{\bm{x}}||_{\mathbb{E}}=D_{\overline{\bm{0}}_{max}}=9.1$. Vectors outside this radius lead to instability and performance degradation due to inaccurate approximation. This problem is only exacerbated as the dimensionality of the hyperbolic vector increases. Higher dimensional vectors tend to have larger norms which limits hyperbolic models’ abilities to scale up.
To constrain hyperbolic vectors within a specified maximum distance, either a normalization function or a parameter clipping method is required. Parameter clipping can be challenging as it may lead to information loss and introduce non-smooth gradients. On the other hand, common normalization functions like tanh and the sigmoid function tend to saturate quickly, limiting their effectiveness as seen in the sigmoid implementation by Chen et al. [6].
@@
To address these issues, we introduce a modified scaling function, designed to provide finer control over both the maximum values and the slope of the curve. A visualization of this function is provided in Figure˜2, and the formulation is presented below:
$\bm{y}_{\text{rescaled}}=\frac{\bm{y}}{\lVert\bm{y}\rVert}\cdot m\cdot\text{tanh}\left(\lVert\bm{y}\rVert\cdot\frac{\text{atanh}(0.99)}{s\cdot m}\right)$ (2)
where $\bm{y}\in\mathbb{R}^{d}$, $m$ is our desired maximum value, and $s$ controls the slope of the curve. We now have a maximum distance value to adhere to and a flexible normalizing function. To apply this to the hyperbolic embeddings, we suggest performing the scaling on the tangent plane of the origin. However, this is an expensive operation to perform repeatedly, as such we derive in Appendix˜C the equivalent factorized form for the scaling of the space values:
$\bm{x}_{s_{\text{rescaled}}}=\bm{x}_{s}\times\frac{e^{\frac{D(\bm{x},\overline{\bm{0}})_{\text{rescaled}}^{K}}{\sqrt{K}}}-e^{\frac{-D(\bm{x},\overline{\bm{0}})_{\text{rescaled}}^{K}}{\sqrt{K}}}}{e^{\frac{D(\bm{x},\overline{\bm{0}})^{K}}{\sqrt{K}}}-e^{\frac{-D(\bm{x},\overline{\bm{0}})^{K}}{\sqrt{K}}}}$} (3)
where $D(\bm{x},\overline{\bm{0}})_{\text{rescaled}}^{K}$ are the distances obtained by plugging $D(\bm{x},\overline{\bm{0}})$ and $D^{K}_{\overline{\bm{0}}_{max}}$ in Equation˜2.
4 Experiments
In order to empirically prove the effectiveness of our proposed solutions, we apply them to metric learning to test low precision learning, EEG classification for generalization ability in data hungry scenarios, and image classification and generation tasks for component efficiency. We also apply the curvature learning for most scenarios to study the new optimizer scheme and possible benefits. We include more detailed ablations and experiments on graph embeddings in the Appendix E.
@@ 4.1 Hierarchical Metric Learning Problem
In their paper Kim et al. [19] take on the problem of hierarchical clustering using an unsupervised hyperbolic loss regularizer they name HIER. This method relies on the use of hierarchical proxies as learnable ancestors of the embedded data points in hyperbolic space. In the following experiment, we extend HIER to the Lorentz model (LHIER) and compare against the results provided by Kim et al. [19].
@@
To adapt the HIER model to the hyperboloid, we first replace the Euclidean layer norm and linear layer with a Lorentzian norm and linear layer [3]. We then modify the HIER loss by replacing the Poincaré distance with the Lorentzian distance and optimizing Lorentzian hierarchical proxy parameters directly on the manifold. We use our new optimization schema and apply distance rescaling before layer norm and after the linear layer.
@@
Kim et al. [19] employ the Euclidean AdamW optimizer with Euclidean parameterizations of hyperbolic proxies. Their experiments use FP16 precision. As the setting is already hyperbolic, significant gains from transitioning to Lorentz space are unlikely. We use this setup to evaluate our components’ ability to learn curvature in low-precision, unstable environments. Any improvements likely stem from better curvature adaptation to the data and task. We follow the experimental setup in Kim et al. [19].
@@ Results
Table 1: Performance of metric learning methods on the four datasets as provided by [19]. $\dagger$ indicates models using larger input images. Network architectures are abbreviated as, R–ResNet50 [41].
@@
NSoftmax [45] R512 61.3 73.9 84.2 90.4 78.2 90.6
ProxyNCA++ [39] R512 69.0 79.8 86.5 92.5 80.7 92.0
Hyp [9] R512 65.5 76.2 81.9 88.8 79.9 91.5
HIER [19] R512 70.1 79.4 88.2 93.0 80.2 91.5
@@
Increase in % 4.71 3.78 2.04 1.08 1.49 1.20
As shown in Table˜1, LHIER learns curvature without stability issues. Our approach improves model performance, with recall@1 gains ranging from 1.49% to 4.71%. Additional ablation experiments with different optimizers and configurations are detailed in Table˜7.
@@ 4.2 EEG Classification Problem
Given EEG recordings labeled into distinct categories, the goal is to classify new recordings based on signal patterns. Each recording is a time series $X\in\mathbb{R}^{C\times T}$, where $C$ is the number of channels and $T$ is the time length. Each label is $y\in\{1,\dots,K\}$. Given $N$ labeled recordings $\left((X_{1},y_{1}),\dots,(X_{N},y_{N})\right)$ from an unknown distribution $p$, the task is to train a model $\hat{y}$ that maps EEG signals $X$ to their correct class.
@@
In their work, Pan et al. [32] introduces Matt, a Riemannian model based on the SPD manifold. The model processes data through convolutional denoisers, extracts embedding covariances, and embeds them using a Riemannian attention mechanism before projecting the outputs back into Euclidean space for classification via a linear layer. Our proposed model, HyperMatt, extends this framework by projecting the resulting covariances onto the hyperboloid instead and processing them with a hyperbolic attention layer [6]. The outputs are then classified directly on the manifold using the MLR [3].
@@
Hyperbolic spaces are prone to overfitting, especially in EEG classification, where models are trained per subject with limited sessions, leading to data scarcity. The proposed RAdamW optimizer theoretically improves regularization, ensuring smoother convergence. We apply our method to the datasets and the evaluation criteria in Pan et al. [32] to test this.
Results
In Table˜2 we show the performance of HyperMAtt compared to the current state-of-the-art baselines, where the results were aggregated from [32, 5]. The performance for HyperMAtt is measured across 10 runs and we report the mean and standard deviation. Here, our new optimizer achieves state-of-the-art results for SSVEP and ERN, the two comparatively smaller datasets, which are more prone to overfitting. We also show the improved performance of our RAdamW optimizer vs the existing RAdam by training the model using both and comparing the results in Table˜8. We also achieve these results in significantly less time/epoch compared to MAtt. Specifically for MI 0.077s/epoch vs 3.4s/epoch, for SSVEP 0.081s/epoch vs 4s/epoch, and for ERN 0.061s/epoch vs 1.9s/epoch, resulting in an average speed-up by a factor of 42.
Models MI SSVEP ERN
[…]
Increase in % -0.78 4.01 2.59
4.3 Standard Image Classification Problem
Problem Setting and Model
In their work, Bdeir et al. [3] propose a fully hyperbolic and hybrid encoder RessNet. To adapt our methods for both models, we introduce the distance rescaling function after every convolution in the encoder. Additionally, we attempt to improve the performance by performing a parametrization trick on existing CUDA convolutions to ensure hyperbolic outputs. This is done by parametrizing the convolution weight as a rotation operation before passing the input, and then applying a boost operation afterwards. The end result is then equivalent to the convolution operation by Bdeir et al. [3] while mitigating computational complexity. Specific implementation details can be found in Appendix˜B. We denote our Hybrid and Fully hyperbolic models as HECNN+ and HCNN+ respectively.
Experimental Goals
In their paper, [3] notice lower performance in the fully hyperbolic models compared to the hybrid models and attribute this to instability in training. Additionally, they cite crashes and model divergence when learning the curvature instead of setting it to constant. This would then be an ideal setting to test the optimization schema and rescaling function and their impact on performance and stability.
Results
In the ResNet-50 experiments in Table˜3, HECNN+ significantly outperforms both the Euclidean model and the base hybrid model across both datasets, demonstrating the positiv effect of curvature learning. We also see a $\sim 48\%$ reduction in memory usage and $\sim 66\%$ reduction in runtime. We attribute this improvement to efficient closed-source CUDA convolution operations we can now leverage. One other benefit that we find from learning the curvature is quicker convergence, where the model is able to reach convergence in 130 epochs vs the 200 epochs required by a static curve model. ResNet-18 experiments show similar findings and can be found in Appendix E.
Table 3: Performance and runtime Analysis for ResNet-50 models. We report classification accuracy (%) and the best performance is highlighted in bold (higher is better). All experiments were conducted on a single NVIDIA RTX 4090 GPU and an AMD EPYC 7543 CPU.
CIFAR-100 Tiny-ImageNet VRAM tepoch
$(\delta_{rel}=0.23)$ $(\delta_{rel}=0.20)$ For Cifar100 For Cifar100
Euclid $78.52$ $66.23$ 4.5GB 30s
HECNN 79.83{79.83} $66.30$ 15.6GB 300s
HECNN+ $\bm{80.86}$ $\bm{67.18}$ 8.1GB 100s
4.4 VAE Image Generation